OpenStudy (anonymous):

Solve. x2 - 7x = -12

OpenStudy (anonymous):

x=4,3

OpenStudy (amoodarya):

\[x^2-7x+12=0\\(x-3)(x-4)=0\\or \\ax^2+bx+c=0\rightarrow x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

1st